OpenStudy (anonymous):

Use multiplication,division,addition and at least one set of parenthesis to write an expression that simplifies to 7, 13 or 17. Demonstrate the consequences of not using proper order of operations by showing that other orders yield different answers.

OpenStudy (anonymous):

can anyone help with this?

OpenStudy (paxpolaris):

Just create an expression like: \[(8+4)\times3\div2+1=19\] Then show that doing the operations in different orders would give you different (incorrect) answers

OpenStudy (turingtest):

\[2\times5-3=(2\times5)-3=10-3=7\]but\[2\times5-3\neq2\times(5-3)=2\times2=4\neq7\]